An object moving at a constant speed of 100 meters in 25 sec.
What is the speed of the object?
How far doed the object move in 20 sec?
How longwill it take the object to move 80m?
1
Expert's answer
2012-09-13T11:36:49-0400
V = S / t = 100 / 25 = 4 m/s is the speed of the object.
In 20 sec: S(20) = V
* 20 = 4 * 20 = 80 m.
To move 80 m: t(80) = 80 / V = 80 / 4 = 20 s.